Can you help with this question?
Say that the variance of a set of data is 24/5. At a minimum, about how many data points will be within 4.3818 of the mean of measurements?
A. 4/5
B. 3/4
C. 2/3
D.1/2

Answer by venugopalramana(3286)

VARIANCE =24/5
RMS DEVIATION =SIGMA=S SAY = SQRT(24/5)=2.191
HENCE THE DEVATION OF 4.3818 IS AMOUNTING TO 2*S VARIATION...AS PER NORMAL DISTRIBUTION OVER 95% OF DATA ARE ON THIS 2S DEVIATION LEVEL.HENCE A IS MOST APPROPRIATE ANSWER THOUGH IT IS ONLY 80%